Prelab questions 1. What is the significance of the statement qsystem + Surroundings zero Law of Thermodynamics? statement
Answer #1

1) This means that entropy is a conserved measure, that is, what is produced by entropy in a system is obtained from another side, in this case from the surroundings.

2) The reaction that occurs is:

NaOH + HCl = NaCl + H2O

Both reagents are limits because they have the same amount of moles and the same stoichiometric coefficient, the moles are:

n = M * V = 1 M * 0.025 L = 0.025 mol

Enthalpy change is calculated:

ΔH = - q / n = 1,295 kJ / 0.025 = 51.8 kJ / mol

3) The NIE that occurs is:

H + + OH- = H2O

The enthalpy change is:

ΔH = ΔHf products - ΔHf reagents = - 285.83 - 0 + 230 = - 55.83 kJ / mol

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.